40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 26 17:07:41 2019, gautam, Update, LSC, POX / POY calibration POX_POY_sensorNoise.pdfALSnoise_20191125.pdf
    Reply  Tue Nov 26 18:16:08 2019, rana, Update, LSC, POX / POY calibration 
       Reply  Wed Nov 27 12:16:52 2019, gautam, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power ITMXshadowSensors.pdfITMYshadowSensors.pdf
          Reply  Wed Nov 27 13:14:02 2019, rana, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power 
             Reply  Wed Nov 27 16:10:29 2019, gautam, Update, LSC, AOM reconnected 
Message ID: 15049     Entry time: Tue Nov 26 17:07:41 2019     Reply to this: 15050
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX / POY calibration 

Summary:

Since we are using the POX and POY photodiodes as out-of-loop sensors for measuring the ALS noise, I decided to double-check their calibrations. I determined the following numbers (for the single arm lock):

POX_I [with 30dB whitening gain]: (8 +/- 1)e-13 m/ct

POY_I [with 18dB whitening gain]: (0.9 +/- 0.1)e-13 m/ct

With this calibration, I measured the in-loop spectra of the XARM and YARM error-points when they are locked - they line up well, see Attachment #1. Note that these numbers are close to what we determined some time ago using the same method (I drove the ITMs then, but yesterday I drove the ETMs, so maybe the more accurate measure of uncertainty is the difference between the two measurements).

Attachment #2 shows the out-of-loop spectra sensed by these photodiodes with this calibration applied, when the arms are under control using ALS beat frequencies as the error signals, and controlled in the CARM/DARM basis. Need to think about why there is such a difference between the two signals.

Methodology:

The procedure used was the same as that outlined here.

  • I started by calibrating the AS55_Q output with the free-swinging Michelson.
  • Next, I lock the Michelson and calibrate the BS and ITM actuators using the newly calibrated AS55_Q.
  • Next, I calibrate the ETM actuator gains by measuring the ratio of response in POX/POY of driving the (unknown) ETMs and the (known) ITMs.
  • Finally, I calibrate the POX/POY photodiodes by driving the ETMs by a known amount of meters (at ~310 Hz where the loop gain is negligible because of the sensing matrix measurement notches).

Summary of DC actuator gains:

Optic Series resistance [ohms] x3 Analog gain? x3 Digital gain? DC gain [nm/ct]
BS 100 No Yes 9.48 +/- 0.01
ITMX 400 No Yes 2.42 +/- 0.01
ITMY 400 No Yes 2.41 +/- 0.01
ETMX 2.2k Yes No 1.23 +/- 0.02
ETMY 400 Yes No 6.62 +/- 0.12

The quoted values of the DC gain are for counts seen at the output of the LSC filter bank. I've attempted to show that once we account for the different series resistance and some extra gains between the output of the LSC filter bank and the actual coil, things are fairly consistent.

Some remarks:

  • I do not understand why we need an extra 12dB of whitening gain on the POX channel to get similar PDH fringe height as the POY channel. The light level on these photodiodes is the same, and the RF transimpedances at 11 MHz are also close according to the wiki (3kohm for POX, 2kohm for POY).
  • At night-time, the ALS noise did indeed get reduced compared to what I measured earlier in the evening.
  • Even assuming 50% error in the calibration factors, it's hard to explain the swing of TRX/TRY when the CARM offset is brought to zero.
  • The increase in (admittedly in-loop) CARM noise as the offset is reduced still seems to me to be correlated with the buildup of IR power in the arm cavities.
Attachment 1: POX_POY_sensorNoise.pdf  62 kB  | Hide | Hide all
POX_POY_sensorNoise.pdf
Attachment 2: ALSnoise_20191125.pdf  35 kB  | Hide | Hide all
ALSnoise_20191125.pdf
ELOG V3.1.3-